OpenStudy (anonymous):

Simplify the expression. 3^7/3^5 A. 335 B. 312 C. 1/3^9 D. 9

OpenStudy (anonymous):

d.9

OpenStudy (diyadiya):

\[\Huge \frac{x^m}{x^n}= x^{m-n}\]

OpenStudy (anonymous):

the rule of exponents is that when you divide them, and the base is the same, then you just subtract the exponents, like @Diyadiya just showed.

OpenStudy (diyadiya):

\[\Large \frac{3^7}{3^5}= 3^{7-5}\] :)

OpenStudy (anonymous):

3^7=2187 3^5=243 2187/243=9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help with another one?

OpenStudy (anonymous):

sure,